### v1

# Performance Gap in Entity Knowledge Extraction Across Modalities in Vision Language Models

Abstract
Vision-language models (VLMs) excel at extracting and reasoning about information from images. Yet, their capacity to leverage internal knowledge about specific entities remains underexplored. This work investigates the disparity in model performance when answering factual questions about an entity described in text versus depicted in an image. Our results reveal a significant accuracy drop—averaging 19%—when the entity is presented visually instead of textually. We hypothesize that this decline arises from limitations in how information flows from image tokens to query tokens. We use mechanistic interpretability tools to reveal that, although image tokens are preprocessed by the vision encoder, meaningful information flow from these tokens occurs only in the much deeper layers. Furthermore, critical image processing happens in the language model’s middle layers, allowing few layers for consecutive reasoning, highlighting a potential inefficiency in how the model utilizes its layers for reasoning. These insights shed light on the internal mechanics of VLMs and offer pathways for enhancing their reasoning capabilities.
Performance Gap in Entity Knowledge Extraction
Across Modalities in Vision Language Models
Figure 1: Illustration of our motivating observation. The model successfully extracts its knowledge about the entity when the entity is represented in text. It further successfully identifies the entity in the image. But it fail to combine the two.
1 Introduction
Over the past several years, vision-language models (VLMs) have made remarkable progress in linking visual inputs with natural language descriptions, enabling them to perform tasks such as image recognition Radford et al. (2021); Mirza et al. (2024), video understanding Lin et al. (2024), and Visual Question Answering (VQA) Liu et al. (2023b, a). Much of the existing research evaluating these models focuses on their capacity to understand and describe information presented in the input image: counting objects, assessing complex visual arrangements, and understanding interactions between humans and objects Dang et al. (2024); Basu et al. (2024). While this has demonstrated that VLMs can integrate visual and textual information effectively, it does not fully address a critical question:
how well can VLMs exploit their internal knowledge about entities presented in an image rather than simply described them in text?
This question opens a new angle on VLM evaluation. Instead of probing their grasp of broad scenes, we consider a different scenario: retrieving factual knowledge about visually presented entities. For example, if the subject of an image is a notable figure, say Robbin Williams, we can ask the model "Who is Robbin William’s spouse?", which we refer to as a textual cue. Or we can provide the model with an image of Robbin Williams and ask “Who is the spouse of the subject of this image?”, which we refer to as a visual cue. The second case can be viewed as a two-hop reasoning task. First, the system must recognize the entity visually—identification. Next, it must connect this recognized entity to its stored factual knowledge—extraction. Figure 1 provides an example where the model contains the knowledge to answer the question, but fails to extract it when provided with a visual representation. A detailed explanation about this evaluation appears in section 2.
Our empirical findings reveal a large accuracy drop when relying on visual cues rather than textual cues. This likely stems from the model investing too many layers in identifying the subject visually, leaving fewer layers for the second hop of factual reasoning (Biran et al., 2024). To investigate these dynamics more concretely, we utilize techniques of mechanistic interpretability such as attention knockout (Geva et al., 2023), activation patching (Vig et al., 2020) and vocabulary projection (nostalgebraist, 2020; Geva et al., 2021). They allow us to intervene in the model’s processing mid-inference, altering how and when image information influences the final answer. By tracking the flow of information through the network, we show that entity recognition and knowledge extraction are deeply entangled processes, with identification dominating the representational space longer than what one would expect. Consequently, the model’s internal knowledge is leveraged less efficiently.
The contribution of our work is providing a better understanding of how VLMs process entities and integrate their internal factual knowledge with image information, highlighting a gap between the processing of visual and textual information. Instead of treating entity-level knowledge extraction as a monolithic process, we reveal that it involves at least two interdependent steps. We show that the initial reliance on visual cues can restrict later factual reasoning compared to textual cues. This understanding highlights the need for more multi-modality specific evaluation protocols. Our experiments, which measure the performance gap and apply attention-based hidden state interventions, provide a foundation for future efforts to enhance the models’ ability to draw on their underlying knowledge when presented with visual inputs.
We provide next, a motivational experiment pointing to a gap between textual and visual representations. Then we present our empirical approach to further analyze it. Related work is deferred to Appendix 4. We release our code and data at https://github.com/ido-co/vlm-modality-gap.
2 Motivation
We start by quantifying the model’s performance gap in entity knowledge extraction when presented with visual versus textual representations of entities. This experiment motivates the subsequent experiments which aim to explore and explain this gap. We use Llava-1.5-7B (Liu et al., 2023b) as the VLM in all experiments.
Experiment.
We checked entity recognition by prompting the model to identify the subject in an image and comparing its response to a list of aliases for that entity, aggregated from Wikidata (Vrandečić and Krötzsch, 2014). Using the IMDB-WIKI dataset (Rothe et al., 2015), we sampled 10,000 entity-image pairs and retained the 300 where the model successfully recognized the entity by prompting the model with an image and the prompt “Identify by name the subject of this image”. This step is necessary to isolate the knowledge extraction capabilities of the model from its ability to recognize entities in images.
[…]
In the textual input case, the model is presented with the question as is. For the visual input, the model receives an image of the entity along with a reformulated query, such as “Where was the subject in the image born?”. This ensures that the model needs to internally identify the subject depicted in the image before reasoning and generating an answer.
Results.
We measured accuracy per entity as the proportion of correctly answered questions for that entity and calculated the overall mean accuracy across all entities. The mean accuracy for textual inputs is 52%, compared to 38% for visual inputs. The average per-entity accuracy drop is 19.5%.
Interestingly, a common mistake made by the model when asked to name the subject’s spouse/parent/sibling/child, is to name the subject itself. This observation leads to our the hypothesis that by the time information about the entity propagates from the image tokens to later token positions, there is not enough computational capacity left to use this information, and the model outputs the entity’s name.
Some entities did exhibit an increase in accuracy with visual inputs. This specifically occurred when the textual input alone was insufficient for reasoning, but the visual input provided contextually inferable details. For instance, a subject wearing a French national team uniform implicitly informed the model that the language spoken by the subject is French.
3 Experiments
We aim to understand at what point in the inference pass the entity in an image is identified, and this identity is propagated to the query and generated token positions. To this end, we will apply interventions in the model’s inference. We continue to focus on entity-image pairs for which, without intervention, the model correctly answers the prompt "Identify by name the subject of this image."
Notation.
Following Liu et al. (2023b), denote by $\mathbf{X}_{\mathbf{v}}$ an input image, and by $\mathbf{H}_{\mathbf{v}}=\mathbf{W}\cdot g(\mathbf{X}_{\mathbf{v}})$ its visual features extracted by the CLIP vision encoder and projected to a sequence of language embedding tokens.
For a text query, $\mathbf{X}_{\mathbf{q}}$, let $\mathbf{H}_{\mathbf{q}}$ denote the sequence of language embedding tokens produced by applying the language tokenizer to $\mathbf{X}_{\mathbf{q}}$.
The input to the language model, with hidden dimension $d$, is formed by concatenating the token sequence $\mathbf{H}_{\mathbf{v}}$ with $\mathbf{H}_{\mathbf{q}}$. Let $\mathbf{h}_{i,v}^{\ell}$ and $\mathbf{h}_{i,q}^{\ell}$ be the hidden representation for the $i^{\text{th}}$ vision and query language tokens. As the language model processes the entire input at each layer $l$, it builds a representation $\left(\mathbf{h}_{1,v}^{\ell}\ldots\mathbf{h}_{n,v}^{\ell},\mathbf{h}_{1,q}^{\ell}\dots\mathbf{h}_{m,q}^{\ell},\mathbf{h}_{1,g}^{\ell}\dots\mathbf{h}_{k,g}^{\ell}\right)$, where $\mathbf{h}_{k,g}^{\ell}$ are the generated tokens hidden representations, $\left|\mathbf{H}_{\mathbf{v}}\right|=n$, $\left|\mathbf{H}_{\mathbf{q}}\right|=m$ and $\ell=0$ refers to the initial token embeddings inputted to the LLM.
In our experiments we focus on how information flows from the vision hidden states $\left(\mathbf{h}_{1,v}^{\ell}\ldots\mathbf{h}_{n,v}^{\ell}\right)$ to the query and generated tokens. We do this by applying techniques of mechanistic interpretability.
Experiments Overview.
In the experiments, we begin in Section 3.1 by locating an upper bound on where the model extracts the identity from the image, using activation patching (Vig et al., 2020) between different entities. Having established an upper bound, we ask if the layers up to this bound are efficiently utilized, or if the model could have extracted the identity sooner. We study this in Section 3.2 using forward activation patching. We defer to the appendix an additional experiment which uses attention knockouts (Geva et al., 2023) and complements the results of Section 3.1.
3.1 Late information flow
We begin by identifying the point of critical information flow from the the image hidden representation to later positions. This will provide an upper bound on the layer depth at which entity identification happens and point us to where this information is likely to propagate onward. We can subsequently infer that by this point the model mostly finalizes its image-based representation enrichment and focuses on leveraging this information for prediction. We hypothesize that this happens in deep layers, leaving few layers for factual knowledge extraction, which may begin to explain the performance gap presented in Section 2.
The experiment involves two distinct forward passes through the model, using different entities: an original entity and a injected entity. Each forward pass is initiated with a prompt asking the model to identify the subject of an image. We extract the image-token representations $\mathbf{h}_{1,v}^{\ell}\ldots\mathbf{h}_{n,v}^{\ell}$ from the injected entity forward pass at a specific layer. These representations are then patched at the same token positions and same layer of the forward pass of the original entity. After patching, we allow the model to complete its forward pass and predict which entity is identified. The layer at which the predicted entity switches from being the injected entity to the original entity identifies the critical point where the model has completed sufficient processing of visual cues such that patching does not affect its prediction.
Results.
Figure 2 shows the effect of patching image tokens across different layers.
Layers 0–19: When image tokens are patched at or before layer 19, the model predominantly predicts the injected entity. This strengthens the hypothesis that the middle layers play a critical role in how the model interprets the image. Our results suggest that despite having 18 layers to process and propagate information about the original entity to later positions, the information conveyed within the narrow window of layers 19 to early twenties ultimately determines the outcome.
Layers 20+: Patching beyond layer 19 results in the model reliably predicting the target entity, suggesting that image-based entity identification is finalized by this stage.
These results suggest two possible processing paths:
I. Parallel identification and knowledge extraction: Generation and query token representations, can be enriched by the image tokens via the attention mechanism concurrently with entity identification before layer 19, implying that the model incrementally combines these processes over several layers, with more impact to information passed in later layers.
II. Sequential processing: Identification may occur just after layer 19, with the remaining layers allocated for factual reasoning and prediction tasks. This would leave a limited subset of late layers for extracting factual knowledge about the entity, potentially leading to degraded question-answering performance.
Figure 2: Entity Patching: Image token hidden states are extracted from the forward pass of the injected entity’s image and injected into the same layer of the original entity’s forward pass. The x axis is the injection layer. Orange line indicates the percent of times the injected entity was generated. Blue is the percent of times the original entity was generated.
3.2 Long processing is necessary in most cases
Having identified the middle layers as the point where critical information propagates from image hidden states to later positions, and having established an upper bound on where identification happens, we will try to answer which of the above mentioned processing paths is more likely, and also answer whether all these layers are truly required for identification? Could it be that the entity was already identified in earlier layers, but was unable to utilize this in the successive layers?
To answer these questions we apply activation patching of the image hidden representations from an early layer into layer 19, by setting $\left(\mathbf{h}_{1,v}^{19}\ldots\mathbf{h}_{n,v}^{19}\right)=\left(\mathbf{h}_{1,v}^{\ell}\ldots\mathbf{h}_{n,v}^{\ell}\right)$ for values of $\ell$ ranging 0 to 18. If identification indeed happens in early layers, "skipping" all the layers until the point of information propagation should retain most of the identification accuracy.
Results.
Figure 3 shows the results, where every bar shows the percent of correctly identified entities after applying forward patching from the layer indicated by axis x to layer 19. From these results we learn that even when patching the representation from layer 0, 40% of entities are still correctly recognized. This suggests that in these cases the representation provided by the vision encoder, encodes sufficient signals to identify the entity. Up to layer 9 there is very little improvement, indicating that these play a lesser role in the identification process. From layer 10 we see a significant improvement until layer 18 where 80% of the entities are identified. This trend indicates that in the majority of cases, the middle layers (10-19) are critical in the identification process.
With these results in mind, we turn back to our motivating experiment and ask whether the QA accuracy is better for entities with early identification. In this experiment, we split our dataset into entities where forward patching from early $(<10)$ layers maintains correct identification, lets call these early-id, and those requiring late layers, late-id. The average accuracy per entity of the early-id entities is 40% compared to 37% for the late-id. This small difference means that the model does not benefit from early identification and this critical information waits for the middle layers to be attended to. This strengthens the case for the second processing pass suggested before, of sequential processing of identification and extraction.
Figure 3: Each bar shows the percent of correctly identified entities after applying forward patching from the layer indicated by axis x to layer 19.
4 Related work
VLMs have emerged as a powerful tool, demonstrating impressive capabilities in tasks such as visual reasoning, object recognition, and multimodal understanding Basu et al. (2024). However, their internal mechanisms and interpretability remain active areas of research. Below, we review the related works categorized into general VLM interpretability, cross-modal information flow, attention mechanisms, and mechanistic approaches to visual token processing.
Recent surveys on large multimodal models provide a comprehensive overview of explainability challenges in these systems. Dang et al. (2024) propose a framework categorizing interpretability across data, model, and training perspectives. They highlight the alignment and decomposition of modalities as key challenges, emphasizing the need for explainability to ensure trustworthiness in multimodal AI applications. In this context, techniques for token-level, neuron-level, and embedding-level interpretability have been systematically analyzed, offering insights into the mechanisms driving decision-making processes.
Previous research has examined the internal information distribution within these models Basu et al. (2024), highlighted the limitations arising from pre-trained vision encoders Tong et al. (2024), and investigated the issue of hallucinations in VLMs Bai et al. (2024). Studies such as Huo et al. (2024) delve into neuron-level interpretations, uncovering how specific neurons contribute to multimodal reasoning. Similarly, Jiang et al. (2024) address hallucination issues in VLMs by examining the influence of vision-language representations. Both works highlight the importance of dissecting model components to improve transparency and robustness.
On the methodological front, attention-based mechanisms remain a cornerstone for interpretability. Tao et al. (2024) explore global and local semantic representations, while Liu et al. (2023b) focus on visual instruction tuning to enhance alignment in cross-modal embeddings. Chefer et al. (2021) provides a technique to highlight the input pixels in an image that affects the decision of a transformer. Chen et al. (2022) further extends this approach for CLIP and Yellinek et al. (2023) suggests using a tree structure to improve the use of HilaCAM Chefer et al. (2021) in VLMs.
Finally, causal approaches and model manipulation techniques have gained traction for evaluating VLMs. Palit et al. (2023) introduce causal tracing tools to understand information flow in vision-language systems, while Zhang et al. (2024) analyze information redundancy and relevance across reasoning tasks. These works demonstrate the efficacy of manipulating hidden states and attention mechanisms to investigate and improve model performance.
Our work is most related to Neo et al. (2024); Kaduri et al. (2024) who investigated processing of information in VLMs. Neo et al. (2024) provided insights into the role of attention layers in capturing visual attributes and object-specific details. Contemporary to us, Kaduri et al. (2024) demonstrated that query tokens store high-level image descriptors. Their findings reveal that middle layers of the model are critical for cross-modal knowledge transfer, while fine-grained visual details are extracted in a spatially localized manner. Yet, both of these works do not study the gap we reveal between the processing of textual and visual representations.
5 Conclusion
This work explores how VLMs process visual entities, enhancing understanding of multi-modal architectures. Our findings highlight that reasoning about entities shown in an image is not a straightforward extension of reasoning about entities described in text, rather, it more closely resembles the processing of multi-hop questions. The model dedicates a substantial portion of its computational "budget" to entity identification and information processing, leaving few layers for the subsequent factual reasoning step. Thus, critical image-related processing often concludes relatively late in the processing flow, limiting the model’s capacity to extract factual knowledge in the remaining layers thus limiting its overall performance. We also find that this issue persists even for early-id entities, that could be recognized earlier, indicating that current models may not be using their representational capacity efficiently.
By highlighting these challenges, our results offer pathways for improvement. Potential directions for tackling the root of this issue include adjusting training regimes to better separate or balance identification and knowledge extraction steps, as well as designing architectural enhancements that reduce entanglement between these stages, which could lead to VLMs that better, and more reliably, utilize their internal knowledge. For instance, this behavior may be mitigated to some extent by using early-fusion models, such as Chameleon. We leave this exploration for future work.
Limitations
The scope of this work focuses on understanding the cause of the observed behavior, and leaves the solutions as future works.
We focused on the performance of Llava-1.5-7B, as it is a canonical architecture in the field, and allows applying interventions with ease. While Llava is a widely used VLM that has been found useful in many applications, newer, early fusion Team (2024) models may exhibit different behavior to the kind described in this work. We believe that studying the gap we report in this work for Llava also for such early fusion models would be an interesting direction for a future research.
When using visual representations, the image may contain features that can steer the model towards or away from the correct answer. For example, a football fan wearing his favorite team’s shirt could be mistaken for a football player. This work does not go into disentangling the effects of such confounders on the observed gap.
Lastly, while our experiments show clear trends in terms of model performance and internal behaviors, they were conducted on a set of images of 300 popular entities. Extending our analysis to a larger set of entities with different properties, such as popularity and ethnicity, is an important direction for future work.
Acknowledgments
We thank Omri Kaduri for his input on Llava mechanics. IC is supported by the Israeli Science Foundation (grant number 1186/18). DG is supported by the Israeli Science Foundation (grant number 1083/24).
Appendix A Additional experiment: Critical information is processed late
To investigate the stages of the model’s processing pipeline where entity identification occurs, we ask the model to identify the subject in an image and employ an attention knockout strategy following the methodology of Geva et al. (2023). This involves selectively prohibiting attention from textual tokens to image tokens at various layers of the model, effectively blocking cross-modal information flow. We do this by inserting $-inf$ in the attention mask at $[x,y]$ where $x$ are the indices of $\mathbf{h}_{1,q}^{\ell}\dots\mathbf{h}_{m,q}^{\ell},\mathbf{h}_{1,g}^{\ell}\dots\mathbf{h}_{k,g}^{\ell}$, $y$ are the indices of $\mathbf{h}_{1,v}^{\ell}\ldots\mathbf{h}_{n,v}^{\ell}$, and $\ell$ is any layer we want to knock out. We conduct this experiment in two directions:
1. Top-down: Knocking out attention from a start layer until the final layer of the model.
2. Bottom-up: Knocking out attention from the model’s first layer until an end layer.
Notably, part of this experiment has been done in concurrent work by (Kaduri et al., 2024), but on different data, and asking different questions. This experiment checks whether the VLM trend follows Geva et al. (2023) which analyze the information flow for factual recall given textual inputs.
Results. Figure 4 presents both types of knockout:
1. Top-down: Knocking out the second half or more of the model’s attention results in complete collapse of model prediction. The sharp change in layers 15-20 indicates that these layers are critical for information flow from the image tokens to the query and generation tokens. Knocking out the last third of layers has a minimal effect on the model’s output, indicating the model no longer attends to the visual tokens after the critical mid-layer processing is complete, demonstrating a transition to text-dominated reasoning.
2. Bottom-up: Shows that unlike the top layers, restricting the attention in early layers has an immediate and progressive effect on the prediction, underscoring their importance. The effect increases around the middle layers until again collapsing the model’s prediction. This shows that there could be some redundancy in the first layers, and the model can begin attending to image tokens quite late in the processing pipeline and still achieve mostly correct entity identification.
These trends mostly agree with Geva et al. (2023) that knock out attention edges of subject tokens in LLMs. This may indicate that the model treats the image tokens similarly to subject textual tokens. One must remember that unlike the textual tokens, the image tokens are already processed and enriched within the vision encoder. This means that although one could hope that image tokens would enjoy extraction at much earlier stages, allowing for contextualization by the text tokens through more layers, the model exhibits suboptimal integration of image information, relying heavily on mid-layer processing while under utilizing earlier layers for cross-modal fusion.
Figure 4: Attention Knockout: query and generated tokens are restricted from attending to the image tokens in all layers from layer 0 until the marker (in orange) or from the marker until the last layer (in blue)
Qualitative results. To complement the quantitative results, we examined qualitative examples of this degradation, in bottom-up knockout. For instance, Table 1 shows the models output for different values of $\ell$ on an image of LeBron James in action. This example aligns with the general observed quantitative trend, showing that entity identification degrades progressively when employing bottom up attention knockout, the model continues to extract useful visual features before ultimately failing and hallucinating. Another such example at the end.
Layer Output
10 LeBron James is an American professional basketball player who was born on December 30, 1979, in Akron, Ohio
19 The image features a man wearing a yellow shirt and a red hat, holding a basketball in his hands. He appears to be a basketball player
20 The image features a man wearing a yellow shirt, standing in front of a crowd of people.
21 The image features a man wearing a suit and tie, standing in front of a building.
22 The image features a man named LeBron James, who is an American professional basketball player.
29 The image features a man named "Mike" who is standing in a field. He is wearing a black shirt
Table 1: An example of bottom up knockout on an image of LeBron James. The layer numbers indicate the final knockout layer. Only layers with meaningful change in output were included for brevity. Note that the model briefly recovers accurate entity identification.
Appendix B Additional Qualitative Examples of Attention Knockout
Bellow is another result of bottom up attention knockout. Every line shows the output of the model after knocking out another layer.
  1. 1.
    Donald Trump is the 45th President of the United States, serving from January 2017 to January 2021.
  2. 2.
    Donald Trump is an American businessman, politician, and television personality. He was born on June 14, 1946
  3. 3.
    Donald Trump is an American businessman, politician, and television personality. He was born on June 14, 1946
  4. 4.
    Donald Trump is an American businessman, politician, and television personality. He was born on June 14, 1946
  5. 5.
    The image features a man wearing a suit and tie, standing in front of a microphone. He appears to be a public speaker or a politician
  6. 6.
    The image features a man wearing a suit and tie, standing in front of a microphone. He appears to be giving a speech or a presentation
  7. 7.
    The image features a man wearing a suit and tie, standing in front of a microphone. He appears to be giving a speech or a presentation
  8. 8.
    The image features a man wearing a suit and tie, standing in front of a microphone. He appears to be giving a speech or a presentation
  9. 9.
    The image features a man with a receding hairline, wearing a suit and tie, and giving a speech.
  10. 10.
    The image features a man with a distinctive hairstyle, sitting at a podium and giving a speech.
  11. 11.
    The image features a man with a distinctive hairstyle, sitting at a podium and giving a speech.
  12. 12.
    The image features a man with short hair, wearing a suit and tie, and giving a speech. He appears to be a professional orator
  13. 13.
    The image features a man with a distinctive hairstyle, wearing a suit and tie.
  14. 14.
    The image features a man with short hair, wearing a suit and tie, and giving a speech. He appears to be a professional orator
  15. 15.
    The image features a man with a beard, wearing a suit and tie, and giving a speech.
  16. 16.
    The image features a man wearing a suit and tie, standing in front of a microphone. He appears to be a public speaker
  17. 17.
    The image features a man wearing a suit and tie, standing in front of a microphone. He appears to be a public speaker
  18. 18.
    The image features a man wearing a suit and tie, standing in front of a blue background. He appears to be a professional or businessman
  19. 19.
    The image features a man wearing a suit and tie, standing in front of a microphone. He appears to be a professional speaker or an announcer
  20. 20.
    The image features a man wearing a suit and tie, standing in front of a microphone. He appears to be a professional speaker or an announcer
  21. 21.
    The image features a man wearing a suit and tie, standing in front of a building. He appears to be a businessman or a professional
  22. 22.
    The image features a man wearing a suit and tie, standing in front of a building. He appears to be a businessman or a professiona
  23. 23.
    The image features a man standing in front of a building, possibly a church. He is wearing a suit and tie
  24. 24.
    The image features a man standing in front of a building, possibly a church. He is wearing a suit and tie
  25. 25.
    The image features a man standing in a field, possibly a farmer or a worker. He is wearing a hat
  26. 26.
    The image features a man standing in a field, possibly a farmer or a worker in the field. He is wearing a hat
  27. 27.
    The image features a man standing in a field, possibly a farmer or a worker in the field. He is wearing a hat
  28. 28.
    The image features a man standing in a field, possibly a farmer or a worker in the field. He is wearing a hat
  29. 29.
    The image features a man standing in a field, possibly a farmer or a worker in the field. He is wearing a hat
  30. 30.
    The image features a man named "Mike" who is wearing a suit and tie. He appears to be a businessman
  31. 31.
    The image features a man named "Mike" who is wearing a suit and tie. He appears to be a businessman
  32. 32.
    The image features a man named "Mike" who is wearing a suit and tie. He appears to be a businessman
Appendix C Resources
All the experiments were conducted using a single A100 80GB or H100 80GB GPU.